Assuming a SUTA tax rate of 4.2% and a SUTA wage base of $11,000, the SUTA tax owed for an employee who has year-to-date earnings prior to the current period of $9,115, earns $2,100 during the current period, and operates in a state with a credit reduction of 1.2% would be $ .
Answer #1
Wages taxable for SUTA = 11000-9115= $1885
SUTA tax owed = 1885*4.2%= $79.17
Note: Credit reduction does not affect the SUTA rate.